OpenStudy (honeymonster):

can you check this for me..i dont think its quite right :S i'll post it in replies

OpenStudy (honeymonster):

\[3x+2+7x-4=-4x\] 3x+7x+4x=4-2 14x=2 \[\frac{ 14x }{ 14 } = \frac{ 2 }{ 14 }\] x=\[\frac{ 1 }{ 7 }\]

OpenStudy (anonymous):

100% Correct

OpenStudy (honeymonster):

ok thank you :)